OpenStudy (anonymous):

Find the equation of the line that passes through the two points (4,−2) and (1,3). Write your answer in standard form.

OpenStudy (azureilai):

slope is \[\frac{ \Delta y }{ \Delta x }\] You can plug in your points to solve for slope What is the standard form? (Different books define them differently, so you will have to tell me which standard from you are using)

OpenStudy (anonymous):

linear i think

OpenStudy (azureilai):

is it y=mx+b?

OpenStudy (anonymous):

yes

OpenStudy (azureilai):

Ok, so you can still use the slope formula I give you to find the slope, then you can use point slope form to find y - intercept \[(y-y _{1})=m(x-x_{1})\] Choose one of your two points to substitute for (x1,y1) and put in your slope to solve for y - intercept

OpenStudy (anonymous):

so what is m=?

OpenStudy (azureilai):

take your two points, find the change in y-value, and divide it over the change in x-value. (refer to my first reply) it is basically \[\frac{ y_{1}-y_{2} }{ x_{1}-x_{2} }\]